Gina is 9 years younger than her sister, Bridgette. Two years ago, Bridgette was twice as old as Gina was. How old is gina now?
Answer by Paul(988)

Her sister's age = x
Her age = x-9
2 years ago her sister's age = x-2
2 years ago her age = x-9-2-->x-11
EQuation:
2(x-11)=x-2
2x-22=x-2
x=20
20-9=11
Hence, Gina's age is 11 and her sister's age is 20.
Paul.